# A Bayesian framework for active object recognition, pose estimation and shape transfer learning through touch

### v1

@@ III-E Particle Filter (PF) for Object Recognition and Pose Estimation
For each object, class $c$, $\boldsymbol{z}^{{}^{\prime}}_{c,k,t}$ will be evaluated based on all previous observations $\boldsymbol{D}_{1:t}$ at locations $\boldsymbol{X}_{1:t}$ to determine its weight $w^{{}^{\prime}}_{c,*,t}$ with respect to the weight of the old MLE particle $\overline{w}^{*}$. In other words,
$\frac{w^{{}^{\prime}}_{c,*,t}}{\overline{w}^{*}}=\frac{p(\boldsymbol{D}_{1:t}|\boldsymbol{z}^{{}^{\prime}}_{c,*,t},\boldsymbol{X}_{1:t})}{p(\boldsymbol{D}_{1:t}|\boldsymbol{z}^{*},\boldsymbol{X}_{1:t})}$ (26)
Combining equation (25) and equation (26),
$w^{{}^{\prime}}_{c,k,t}=\frac{p(\boldsymbol{D}_{s}|\boldsymbol{z}^{{}^{\prime}}_{c,k,t},\boldsymbol{X}_{s})}{p(\boldsymbol{D}_{s}|\boldsymbol{z}^{{}^{\prime}}_{c,*,t},\boldsymbol{X}_{s})}\frac{p(\boldsymbol{D}_{1:t}|\boldsymbol{z}^{{}^{\prime}}_{c,*,t},\boldsymbol{X}_{1:t})}{p(\boldsymbol{D}_{1:t}|\boldsymbol{z}^{*},\boldsymbol{X}_{1:t})}\overline{w}^{*}$ (27)
@@
One key characteristic of this approximation is that $w^{{}^{\prime}}_{c,k,t}$ can be higher than $\overline{w}^{*}$ only if $\boldsymbol{z}^{{}^{\prime}}_{c,k,t}$ is a better fit compared to $\boldsymbol{z}^{*}$ over all observations $D_{1:t}$.
@@
To further reduce the computational cost, the proposed sampling scheme is only performed when the MLE model evidence $p(\boldsymbol{D}_{1:t}|\boldsymbol{z}^{*},\boldsymbol{X}_{1:t})$ smaller than a certain threshold $\lambda$. In other words, if the MLE is a good fit for the data, new particles will not be proposed.
@@
$\boldsymbol{z}_{j,t}$ := class and the pose of particle $j$ at time step $t$
@@
$\boldsymbol{S}\leftarrow{((\boldsymbol{z}_{1,1},{w}_{1,1}),\cdots,(\boldsymbol{z}_{N,1},{w}_{N,1}))}$
@@
       for $(\boldsymbol{z}_{j,t},\overline{w}_{j,t})\in\boldsymbol{S}$ do
@@
      Resampling:
       $S\leftarrow$ Stochastic Universal Sampling(S)
       Tracking model evidence: Find the MLE particle $(\boldsymbol{z}^{*},\overline{w}^{*})$ based on (11) and calculate the model evidence $p(\boldsymbol{D}_{1:t}|\boldsymbol{z}^{*},\boldsymbol{X}_{1:t})$
       if $p(\boldsymbol{D}_{1:t}|\boldsymbol{z}^{*},\boldsymbol{X}_{1:t})\leq\lambda$ then
             Sampling particles: Sample new particles $(\boldsymbol{z}^{{}^{\prime}}_{k,t},w^{{}^{\prime}}_{k,t})$ based on point-pair features of observed point pairs and (25)-(27), in total $K$ particles are sampled.
@@
                   $\boldsymbol{S}\leftarrow\boldsymbol{S}\cup(\boldsymbol{z}^{{}^{\prime}}_{k,t},w^{{}^{\prime}}_{k,t})$
@@
III-F Discriminating Known and Unknown Objects
@@
To discriminate between known and unknown objects, a threshold needs to be defined for the MLE model evidence $p(\boldsymbol{D}|\boldsymbol{z}^{*},\boldsymbol{X})$. In this study, two types of data points are present, namely, contact points and non-contact points. The criterion for an object to be classified as a known object is defined as follows
$p(\boldsymbol{D}|\boldsymbol{z}^{*},\boldsymbol{X})\geq\frac{1}{Z}(0.90)^{n_{pos}}*(0.50)^{n_{neg}},$ (28)
@@
where $n_{pos}$ denotes the number of observed contact points, $n_{neg}$ denotes the number of observed non-contact points. Recall from equation (14) that, when a contact point is exactly on the estimated surface, its likelihood equals $\frac{1}{Z}$. The criterion encourages the average observation likelihood of contact points to be greater than or equal to $90$ percent of the maximum likelihood of exact contact, while the average observation likelihood of the non-contact points to be greater than or equal to $0.5$. If the observation likelihood for a non-contact point is below $0.5$, it penetrates the MLE shape. The $0.5$ likelihood threshold implies that the MLE shape should not be in penetration with the non-contact points for known objects.
III-G Gaussian Process Implicit Surface for Shape Reconstruction
@@
When a novel object is detected, the Gaussian Process Implicit Surface (GPIS)[14] is used to reconstruct the shape given a prior function $\boldsymbol{\mu}$ and observed data $\boldsymbol{D}$. The aim is to learn a signed distance function that fits the data $\boldsymbol{D}$ while taking into account the prior $\boldsymbol{\mu}$. Similar to $\boldsymbol{\widehat{d}}$ in equation (6), $\boldsymbol{\mu}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{4}$ maps a point $\boldsymbol{x}$ in the Euclidean space to a predictive signed distance value and a gradient vector.
@@
Under the GP assumption, given an unexplored location $\boldsymbol{x}^{*}$, observed contact points $\boldsymbol{X}$ and observed tactile data $\boldsymbol{D}$, the predictive observation $\boldsymbol{d}^{*}$ at $\boldsymbol{x}^{*}$ satisfies,
$\boldsymbol{d}^{*}(\boldsymbol{x}^{*})\sim N(\boldsymbol{\mu}^{*}(\boldsymbol{x}^{*}),\boldsymbol{\Sigma}^{*}(\boldsymbol{x}^{*})).$ (29)
The mean $\boldsymbol{\mu}^{*}(\boldsymbol{x}^{*})$ and the covariance matrix $\boldsymbol{\Sigma}^{*}(\boldsymbol{x}^{*})$ are given by
$\boldsymbol{\mu}^{*}(\boldsymbol{x}^{*})$ $=\boldsymbol{\mu}(\boldsymbol{x}^{*})+\boldsymbol{k}_{*}(\boldsymbol{K}+\sigma^{2}\boldsymbol{I})^{-1}(\boldsymbol{\overline{D}}-\overline{\boldsymbol{\mu}}(\boldsymbol{X})),$ (30)
$\boldsymbol{\Sigma}^{*}(\boldsymbol{x}^{*})$ $=\boldsymbol{k}_{**}-\boldsymbol{k}_{*}(\boldsymbol{K}+\sigma^{2}\boldsymbol{I})^{-1}\boldsymbol{k}_{*}^{T}$ (31)
@@
where $\boldsymbol{k}_{**}$, $\boldsymbol{k}_{*}$ and $\boldsymbol{K}$ is the covariance matrix between $(\boldsymbol{d}^{*},\boldsymbol{d}^{*})$, $(\boldsymbol{d}^{*},\boldsymbol{D})$ and $(\boldsymbol{D},\boldsymbol{D})$ respectively; $\sigma$ denotes the sensory noise level; $\boldsymbol{\overline{D}}$ represents a flattened vector that contains observed data $\boldsymbol{D}$ observed at $\boldsymbol{X}$, while $\overline{\boldsymbol{\mu}}(\boldsymbol{X})$ represents a flattened vector that contains the prior predictive observed values at $\boldsymbol{X}$. The covariance matrix between $\boldsymbol{d}$ at $\boldsymbol{x}$ and $\boldsymbol{d}^{{}^{\prime}}$ at $\boldsymbol{x}^{{}^{\prime}}$ can be calculated as
$cov(\boldsymbol{d},\boldsymbol{d^{{}^{\prime}}})=\begin{bmatrix}k_{f}(\boldsymbol{x},\boldsymbol{x^{{}^{\prime}}})&\nabla_{\boldsymbol{x}^{{}^{\prime}}}k_{f}(\boldsymbol{x},\boldsymbol{x^{{}^{\prime}}})\\ \nabla_{\boldsymbol{x}}k_{f}(\boldsymbol{x},\boldsymbol{x^{{}^{\prime}}})&\nabla_{\boldsymbol{x}^{{}^{\prime}}}\nabla_{\boldsymbol{x}}k_{f}(\boldsymbol{x},\boldsymbol{x^{{}^{\prime}}})\\ \end{bmatrix},$ (32)
where $k_{f}$ is the kernel function of choice. Following (32), $\boldsymbol{k}_{**}$, $\boldsymbol{k}_{*}$ and $\boldsymbol{K}$ can be calculated.
@@
$k_{f}(\boldsymbol{x},\boldsymbol{x}^{{}^{\prime}})=k_{TP}(\boldsymbol{x},\boldsymbol{x}^{{}^{\prime}})=a(2d^{3}-3Rd^{2}+R^{3})$ (33)
$d=||\boldsymbol{x}-\boldsymbol{x}^{{}^{\prime}}||_{2}$ (34)
The first and second derivatives of $k_{TP}(\boldsymbol{x},\boldsymbol{x}^{{}^{\prime}})$ can be calculated as
$\frac{\partial{k_{TP}}(\boldsymbol{x},\boldsymbol{x^{{}^{\prime}}})}{\partial{x_{i}}}$ $=6a(x_{i}-x^{{}^{\prime}}_{i})(d-R),$ (35)
$\frac{\partial^{2}{k_{TP}}(\boldsymbol{x},\boldsymbol{x^{{}^{\prime}}})}{\partial{x_{i}}\partial{x^{{}^{\prime}}_{j}}}$ $=-6a\left(\frac{(x_{i}-x^{{}^{\prime}}_{i})(x_{j}-x^{{}^{\prime}}_{j})}{d}+\delta_{ij}(d-R)\right),$ (36)
@@
where $\delta_{ij}$ denotes the Dirac delta function $\delta(i-j)$. Differing from the definition in [14][21], a scaling coefficient $a$ is introduced to allow the GP to adapt to different novel objects. The higher $a$ is, the larger the assumed process noise is, which corresponds to assuming a larger difference between the prior and the actual object. The kernel parameter $a$ can be updated online using a gradient-based optimizer to maximize the data likelihood of the GP[32]. Last but not least, only contact points are used to update the GP since the signed distance values at contact points are precise, while it is not the case for non-contact points.
III-H Combining PF and GPIS
@@
For GPIS reconstruction, given a fixed prior function $\boldsymbol{\mu}$, the negative log-likelihood of observing data $\boldsymbol{D}$ at a set of points $\boldsymbol{X}$ satisfies the following relation:
$-\ln p(\boldsymbol{D}|\boldsymbol{X})\propto\left(\overline{\boldsymbol{\mu}}(\boldsymbol{X})-\overline{\boldsymbol{D}}\right)^{T}$ (37)
$\boldsymbol{K}^{-1}\left(\overline{\boldsymbol{\mu}}(\boldsymbol{X})-\overline{\boldsymbol{D}}\right)$
@@
Assuming the prior function for the GP is determined by the latent variables $\boldsymbol{z}$, one can choose the prior function by minimizing the log-likelihood function,
$\operatorname*{argmin}_{\boldsymbol{z}}-\ln p(\boldsymbol{D},\boldsymbol{z}|\boldsymbol{X})\propto\left(\overline{\widehat{\boldsymbol{D}}}(\boldsymbol{z},\boldsymbol{X})-\overline{\boldsymbol{D}}\right)^{T}$ (38)
$\boldsymbol{K}^{-1}\left(\overline{\widehat{\boldsymbol{D}}}(\boldsymbol{z},\boldsymbol{X})-\overline{\boldsymbol{D}}\right),$
where $\overline{\widehat{\boldsymbol{D}}}(\boldsymbol{z},\boldsymbol{X})$ denotes a flattened vector that contains the predictive observed signed distance values and normal vectors at locations $\boldsymbol{X}$. Equation (38) takes a similar form to equation (16). The difference is that the diagonal matrix $\boldsymbol{\Sigma}$ is replaced with the inverse of the kernel matrix $\boldsymbol{K}^{-1}$. The MLE particle from the PF has a straightforward interpretation: the combination of object class and pose that minimizes the total square difference between the prediction and observation for all data points.
@@
In the PF, with the known object assumption, there is no spatial correlation between data points. For the GPIS, the test object is assumed to be unknown, therefore spatial correlation stemming from the GPIS’s smoothness assumption is necessary for inferring the signed distance values and the surface normal vectors at unexplored locations.
@@
The GPIS provides the ability of adapting to, and learning, novel shapes, whereas the PF can find the maximum likelihood combination of known object class and poses based on the observed data. Therefore, we propose to use the MLE particles of the PF as a prior for GPIS shape reconstruction. In this manner, the GPIS has a more flexible, grounded prior selection process and the knowledge from known shapes can be transferred to learn new shapes.
@@
It is worth pointing out that, the MLE prior is not guaranteed to be better than other arbitrary priors given that the real object can differ substantially from any known object, but it will always base it’s choice on it’s prior knowledge and the observed data.
III-I Active Data Acquisition

### v3

@@ III-B Customized Particle Filter for Object Recognition and Pose Estimation
For each object, class $c$, $\mathbf{z}^{{}^{\prime}}_{c,k,t}$ will be evaluated based on all previous observations $\mathbf{D}_{1:t}$ at locations $\mathbf{X}_{1:t}$ to determine its weight $w^{{}^{\prime}}_{c,*,t}$ with respect to the weight of the old MAP particle $\overline{w}^{*}$. In other words,
$\frac{w^{{}^{\prime}}_{c,*,t}}{\overline{w}^{*}}=\frac{p(\mathbf{D}_{1:t}|\mathbf{z}^{{}^{\prime}}_{c,*,t},\mathbf{X}_{1:t})}{p(\mathbf{D}_{1:t}|\mathbf{z}^{*},\mathbf{X}_{1:t})}$ (25)
Combining (24) and (25),
$w^{{}^{\prime}}_{c,k,t}=\frac{p(\mathbf{D}_{s}|\mathbf{z}^{{}^{\prime}}_{c,k,t},\mathbf{X}_{s})}{p(\mathbf{D}_{s}|\mathbf{z}^{{}^{\prime}}_{c,*,t},\mathbf{X}_{s})}\frac{p(\mathbf{D}_{1:t}|\mathbf{z}^{{}^{\prime}}_{c,*,t},\mathbf{X}_{1:t})}{p(\mathbf{D}_{1:t}|\mathbf{z}^{*},\mathbf{X}_{1:t})}\overline{w}^{*}$ (26)
@@
One key characteristic of this approximation is that $w^{{}^{\prime}}_{c,k,t}$ can be higher than $\overline{w}^{*}$ only if $\mathbf{z}^{{}^{\prime}}_{c,k,t}$ is a better fit compared to $\mathbf{z}^{*}$ over all observations $D_{1:t}$.
@@
To further reduce the computational cost, the proposed sampling scheme is only performed when the MAP model evidence $p(\mathbf{D}_{1:t}|\mathbf{z}^{*},\mathbf{X}_{1:t})$ is smaller than a certain threshold $\lambda$. In other words, if the MAP is a good fit for the data, new particles will not be proposed.
@@
  $\mathbf{z}_{j,t}$ := class and the pose of particle $j$ at time step $t$
@@
  $\mathbf{S}\leftarrow{((\mathbf{z}_{1,1},{w}_{1,1}),\cdots,(\mathbf{z}_{N,1},{w}_{N,1}))}$
@@
     for $(\mathbf{z}_{j,t},\overline{w}_{j,t})\in\mathbf{S}$ do
@@
     Resampling: $S\leftarrow$ Stochastic Universal Sampling(S)
     Tracking model evidence: Find the MAP particle $(\mathbf{z}^{*},\overline{w}^{*})$ based on (17) and calculate the model evidence $p(\mathbf{D}_{1:t}|\mathbf{z}^{*},\mathbf{X}_{1:t})$
     if $p(\mathbf{D}_{1:t}|\mathbf{z}^{*},\mathbf{X}_{1:t})\leq\lambda$ then
        Sampling particles: Sample new particles $(\mathbf{z}^{{}^{\prime}}_{k,t},w^{{}^{\prime}}_{k,t})$ based on point-pair features of observed point pairs and (24)-(26), in total $K$ particles are sampled.
@@
           $\mathbf{S}\leftarrow\mathbf{S}\cup(\mathbf{z}^{{}^{\prime}}_{k,t},w^{{}^{\prime}}_{k,t})$
@@
III-C Discriminating between Known and Novel Objects
@@
To discriminate between known and novel objects, a threshold needs to be defined for the MAP model evidence $p(\mathbf{D}|\mathbf{z}^{*},\mathbf{X})$. In this study, two types of data points are present, namely, contact points and non-contact points. The criterion for an object to be classified as a known object is defined as follows
$p(\mathbf{D}|\mathbf{z}^{*},\mathbf{X})\geq\frac{1}{Z}(0.90)^{n_{pos}}*(0.50)^{n_{neg}},$ (27)
@@
where $n_{pos}$ denotes the number of observed contact points, and $n_{neg}$ denotes the number of observed non-contact points. Recall from (12) that, when a contact point is exactly on the estimated surface, its likelihood equals $\frac{1}{Z}$. The criterion encourages the average observation likelihood of contact points to be greater than or equal to $90$ percent of the maximum likelihood of exact contact, while the average observation likelihood of the non-contact points to be greater than or equal to $0.5$. If the observation likelihood for a non-contact point is below $0.5$, it penetrates the MAP shape. The $0.5$ likelihood threshold implies that the MAP shape should not be in penetration with the non-contact points for known objects.
III-D Gaussian Process Implicit Surface for Shape Reconstruction
@@
When a novel object is detected, the Gaussian Process Implicit Surface (GPIS) [11] is used to reconstruct the shape given a prior function $\boldsymbol{\mu}$ and observed data $\mathbf{D}$. The aim is to learn a signed distance function that fits the data $\mathbf{D}$ while taking into account the prior $\boldsymbol{\mu}$. Similar to $\mathbf{\widehat{d}}$ in (7), $\boldsymbol{\mu}:\mathbb{R}^{3}\rightarrow\mathbb{R}^{4}$ maps a point $\mathbf{x}$ in the Euclidean space to a predicted signed distance value and a gradient vector.
@@
Under the GP assumption, given an unexplored location $\mathbf{x}^{*}$, observed contact points $\mathbf{X}$ and observed tactile data $\mathbf{D}$, the predictive observation $\mathbf{d}^{*}$ at $\mathbf{x}^{*}$ satisfies,
$\mathbf{d}^{*}\sim N(\boldsymbol{\mu}_{p}(\mathbf{x}^{*}),\boldsymbol{\sigma}_{p}(\mathbf{x}^{*})).$ (28)
The mean $\boldsymbol{\mu}_{p}(\mathbf{x}^{*})$ and the covariance matrix $\boldsymbol{\sigma}_{p}(\mathbf{x}^{*})$ are given by
$\boldsymbol{\mu}_{p}(\mathbf{x}^{*})$ $=\boldsymbol{\mu}(\mathbf{x}^{*})+\mathbf{k}_{*}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}(\mathbf{\overline{D}}-\overline{\boldsymbol{\mu}}(\mathbf{X})),$ (29)
$\boldsymbol{\sigma}_{p}(\mathbf{x}^{*})$ $=\mathbf{k}_{**}-\mathbf{k}_{*}(\mathbf{K}+\sigma^{2}\mathbf{I})^{-1}\mathbf{k}_{*}^{T}$ (30)
@@
where $\sigma$ denotes the sensory noise level. $\mathbf{\overline{D}}$ represents a flattened vector that contains observed data $\mathbf{D}$ observed at $\mathbf{X}$, while $\overline{\boldsymbol{\mu}}(\mathbf{X})$ represents a flattened vector that contains the prior predictive observed values at $\mathbf{X}$; $\mathbf{k}_{**}$, $\mathbf{k}_{*}$ and $\mathbf{K}$ is the covariance matrix between $(\mathbf{d}^{*},\mathbf{d}^{*})$, $(\mathbf{d}^{*},\mathbf{D})$ and $(\mathbf{D},\mathbf{D})$ respectively. A covariance matrix between observation $\mathbf{d}$ at $\mathbf{x}$ and observation $\mathbf{d}^{{}^{\prime}}$ at $\mathbf{x}^{{}^{\prime}}$ can be calculated as
$cov(\mathbf{d},\mathbf{d^{{}^{\prime}}})=\begin{bmatrix}k_{f}(\mathbf{x},\mathbf{x^{{}^{\prime}}})&\frac{\partial}{\partial\mathbf{x}^{{}^{\prime}}}\left(k_{f}(\mathbf{x},\mathbf{x}^{{}^{\prime}})\right)\\ \frac{\partial^{T}}{\partial\mathbf{x}}\left(k_{f}(\mathbf{x},\mathbf{x^{{}^{\prime}}})\right)&\frac{\partial^{T}}{\partial\mathbf{x}^{{}^{\prime}}}\frac{\partial}{\partial\mathbf{x}}\left(k_{f}(\mathbf{x},\mathbf{x^{{}^{\prime}}})\right)\\ \end{bmatrix},$ (31)
where $k_{f}$ is the kernel function of choice. Following (31), $\mathbf{k}_{**}$, $\mathbf{k}_{*}$ and $\mathbf{K}$ can be calculated.
@@
$k_{f}(\mathbf{x},\mathbf{x}^{{}^{\prime}})=k_{TP}(\mathbf{x},\mathbf{x}^{{}^{\prime}})=a(2d^{3}-3Rd^{2}+R^{3})$ (32)
$d=||\mathbf{x}-\mathbf{x}^{{}^{\prime}}||_{2}$ (33)
The first and second derivatives of $k_{TP}(\mathbf{x},\mathbf{x}^{{}^{\prime}})$ can be calculated as
$\frac{\partial{k_{TP}}(\mathbf{x},\mathbf{x^{{}^{\prime}}})}{\partial{x_{i}}}$ $=6a(x_{i}-x^{{}^{\prime}}_{i})(d-R),$ (34)
$\frac{\partial^{2}{k_{TP}}(\mathbf{x},\mathbf{x^{{}^{\prime}}})}{\partial{x_{i}}\partial{x^{{}^{\prime}}_{j}}}$ $=-6a\left(\frac{(x_{i}-x^{{}^{\prime}}_{i})(x_{j}-x^{{}^{\prime}}_{j})}{d}+\delta_{ij}(d-R)\right),$ (35)
@@
where $\delta_{ij}$ denotes the Dirac delta function. Differing from the definition in [11][18], a scaling coefficient $a$ is introduced to allow the GP to adapt to various novel objects. The higher $a$ is, the larger the assumed process noise is, which corresponds to assuming a larger difference between the prior and the actual object. The kernel parameter $a$ can be updated online using a gradient-based optimizer to maximize the data likelihood of the GP [30]. Last but not least, only contact points are used to update the GP since the signed distance values at contact points are exact, while it is not the case for non-contact points.
III-E Combining PF and GPIS
@@
For GPIS reconstruction, given a fixed prior function $\boldsymbol{\mu}$, the likelihood of observing data $\mathbf{D}$ at a set of points $\mathbf{X}$ satisfies the following relation:
$p(\mathbf{D}|\mathbf{X})\propto\exp\left[\left(\overline{\boldsymbol{\mu}}(\mathbf{X})-\overline{\mathbf{D}}\right)^{T}\mathbf{K}^{-1}\left(\overline{\boldsymbol{\mu}}(\mathbf{X})-\overline{\mathbf{D}}\right)\right]$ (36)
@@
Assuming the prior function for the GP is determined by the latent variables $\mathbf{z}$, one can choose the prior function by maximizing the likelihood function,
$\operatorname*{\text{arg max}}_{\mathbf{z}}\exp\left[\left(\overline{\boldsymbol{\mu}}(\mathbf{z},\mathbf{X})-\overline{\mathbf{D}}\right)^{T}\mathbf{K}^{-1}\left(\overline{\boldsymbol{\mu}}(\mathbf{z},\mathbf{X})-\overline{\mathbf{D}}\right)\right].$ (37)
@@
Notice the likelihood function from (36) takes a similar form to (12). The main difference is that the diagonal matrix $\boldsymbol{\sigma}$ is replaced with the inverse of the kernel matrix $\mathbf{K}^{-1}$. In the PF, with the known object assumption, there is no spatial correlation between data points. For the GPIS, the test object is assumed to be novel, therefore spatial correlation stemming from the GPIS’s smoothness assumption is necessary for inferring the signed distance values and the surface normal vectors at unexplored locations.
@@
The GPIS provides the ability of adapting to, and learning, novel shapes, whereas the PF can find the maximum likelihood combination of known object class and poses based on the observed data. Therefore, we propose to use the MAP particles of the PF as a prior for GPIS shape reconstruction. In this manner, the GPIS has a more flexible, grounded prior selection process and the knowledge from known shapes can be transferred to learn new shapes.
@@
It is worth pointing out that the MAP prior is not guaranteed to be better than other arbitrary priors, given that the real object can differ substantially from any known object, however it will always base its choice on the prior knowledge and the observed data.
III-F Active Data Acquisition